OpenStudy (anonymous):

find the area of the shaded region: r=sqrt(theta) theta between 0 and -3pi/2

OpenStudy (anonymous):

\[-\int\limits_{0}^{3\pi/2}d \theta \int\limits_{0}^{\sqrt(\theta)}rdr\]

OpenStudy (anonymous):

Thanks!

OpenStudy (anonymous):

you're welcome did that work?

OpenStudy (anonymous):

How can I verify it?

OpenStudy (anonymous):

9pi^2/16 do you have the answers?